If you are finding proportions, then if it is a decimal, what would this answer be, and please explain how you got the answer, 67.2/g = 16.8/3.3
Answer by eperette(173)

67.2/g = 16.8/3.3
3.3(67.2/g) = 3.3(16.8/3.3)
221.76/g = 55.44/3.3
221.76/g = 16.8
g(221.76/g)= g(16.8)
221.76g/g= 16.8g
221.76 = 16.8g
221.76/16.8 = 16.8g/16.8
13.2 = 1g
or
g= 13.2
your objecttive is to get the g by itself (notice not 1/g, but g) ...so you go by doing inverse operations (i.e. to get rid of division you multiply, to get rid of multiplication you divide)until you get g by itself